# SeaS: Few-shot Industrial Anomaly Image Generation with Separation and Sharing Fine-tuning

### v1

@@ Abstract
Current segmentation methods typically require many training images and precise masks, while insufficient anomaly images hinder their application in industrial scenarios. To address such an issue, we explore producing diverse anomalies and accurate pixel-wise annotations. By observing the real production lines, we find that anomalies vary randomly in shape and appearance, whereas products hold globally consistent patterns with slight local variations. Such a characteristic inspires us to develop a Separation and Sharing Fine-tuning (SeaS) approach using only a few abnormal and some normal images. Firstly, we propose the Unbalanced Abnormal (UA) Text Prompt tailored to industrial anomaly generation, consisting of one product token and several anomaly tokens. Then, for anomaly images, we propose a Decoupled Anomaly Alignment (DA) loss to bind the attributes of the anomalies to different anomaly tokens. Re-blending such attributes may produce never-seen anomalies, achieving a high diversity of anomalies. For normal images, we propose a Normal-image Alignment (NA) loss to learn the products’ key features that are used to synthesize products with both global consistency and local variations. The two training processes are separated but conducted on a shared U-Net. Finally, SeaS produces high-fidelity annotations for the generated anomalies by fusing discriminative features of U-Net and high-resolution VAE features. The extensive evaluations on the challenging MVTec AD and MVTec 3D AD dataset (RGB images) demonstrate the effectiveness of our approach. For anomaly image generation, on MVTec AD dataset, we achieve 1.88 on IS and 0.34 on IC-LPIPS, while on the MVTec 3D AD dataset, we obtain 1.95 on IS and 0.30 on IC-LPIPS. For the downstream task, by using our generated anomaly image-mask pairs, three common segmentation methods achieve an average 11.17% improvement on IoU on MVTec AD dataset, and a 15.49% enhancement in IoU on the MVTec 3D AD dataset. The source code is available at https://github.com/HUST-SLOW/SeaS.
@@ 1 Introduction
Existing segmentation approaches require a large number of anomaly images with mask annotations, while the scarcity of anomaly images obstructs their application in industrial scenarios. To solve this problem, generative methods for industrial scenarios have emerged to expand the training set of segmentation models.
@@
Figure 1: (a) Anomaly Generation (AG) method only generates anomaly and edits it onto normal images guided by the anomaly masks. (b) Anomaly Image Generation (AIG) methods generate anomalies and the industrial products that they lie in simultaneously. (i) DFMGAN trains a generation model for each anomaly type. (ii) SeaS trains a shared generation model for multiple anomaly types.
To the best of our knowledge, generation approaches (Zavrtanik et al., 2021)in industrial scenarios can be broadly classified into two categories: Anomaly Generation (AG) and Anomaly Image Generation (AIG). AG methods (Li et al., 2021; Zavrtanik et al., 2021; Schlüter et al., 2022; Hu et al., 2024)generate anomalies only and merge them into the real normal images using different strategies, e.g., CutPaste (Li et al., 2021)pastes a cropped normal region to normal images, which simulates anomalies by misalignment. AnomalyDiffusion (Hu et al., 2024)generates anomalies by a diffusion model, and edits anomalies onto the normal images guided by the anomaly masks, as shown in Fig. 1(a). However, AG methods require anomaly masks as inputs, which easily suffer from low fidelity and consistency in generation if these masks are unreasonably positioned. In contrast to AG, as shown in Fig. 1(b), AIG approaches take a step further, generating anomalies and the industrial products that they lie in simultaneously. Therefore, AIG faces greater challenges due to the requirement of high-fidelity of both anomalies and products. Previously, AIG methods (Duan et al., 2023; Zhang et al., 2021a; Niu et al., 2020) are mainly based on Generative Adversarial Network (GAN). However, they are limited by insufficient generalization of GAN, lacking specific control over products/anomalies, and inaccurate masks.
@@
In real industrial manufacturing, the products in an individual production line are almost similar to each other, while the anomalies are unforeseeable in shape and appearance. Such an observation reveals a differentiated characteristic, i.e., the products satisfy global consistency with minor variations in local details, while the anomalies hold randomness, which is rarely discussed in existing AG or AIG approaches. Motivated by such a characteristic, we propose a Separation and Sharing Fine-tuning method, short by SeaS, a controllable AIG method based on Stable Diffusion (Rombach et al., 2022). The key idea is to employ Unbalanced Abnormal (UA) Text Prompts containing a set of tokens that characterize products and anomalies separately, so that the anomaly tokens align with the anomaly semantics for diverse generations, and a product token expresses a globally consistent product surface. Specifically, to learn highly-diverse anomalies, we first propose a Decoupled Anomaly Alignment (DA) loss to bind the attributes of the anomalies to different anomaly tokens. Recombining the decoupled attributes may produce anomalies that have never been seen in the training dataset, therefore increasing the diversity of the generated anomalies. Secondly, to learn globally-consistent patterns from products, we propose the Normal-image Alignment (NA) loss. It enables the network to learn the key features of the product from normal images and fine-tune a learnable embedding. Such an embedding ensures the preservation of global consistency amidst local detail variations. Thirdly, according to the experimental analysis, we find that existing methods leverage the low-resolution features to predict the mask, which may introduce a large amount of boundary uncertainty. Thus, we propose a Refined Mask Prediction (RMP) branch to produce pixel-wise anomaly annotations for other downstream tasks. It combines the discriminative U-Net features and high-resolution VAE features to generate accurate and crisp masks in a progressive way. Extensive experiments on AIG and downstream anomaly segmentation tasks show that SeaS outperforms the existing industrial anomaly generation methods. On MVTec AD dataset, our model achieves 1.88 on IS metric and 0.34 on IC-LPIPS. Furthermore, training on the image-mask pairs generated by SeaS, the downstream segmentation models achieve improvements of average +5.53% AP and +11.17% IoU. On MVTec AD 3D dataset (RGB images), our method attains 1.95 on IS metric and 0.30 on IC-LPIPS. Using the image-mask pairs generated by SeaS to train the downstream segmentation models, we exhibit average improvements of +12.13% AP and +15.49% IoU.
@@
  • We reveal different characteristics of products and anomalies, which motivates us to propose SeaS, a novel AIG method. It independently learns products and anomalies on a shared U-Net and ensures the randomness of anomalies and global consistency of products.
  • We propose a Refined Mask Prediction branch to produce accurate and crisp pixel-wise annotations for generated anomalies, which combines the advantages of the discriminative U-Net features and the high-resolution VAE decoder features.
  • Extensive experiments on anomaly image generation and downstream anomaly segmentation tasks show that SeaS outperforms the existing industrial anomaly generation methods.
@@ 2 Related Work
Anomaly Image Generation. Early non-generative methods (DeVries & Taylor, 2017; Li et al., 2021; Zavrtanik et al., 2021) use data augmentation techniques to create anomaly images. Data augmentation techniques lack consistency in anomaly images, resulting in low fidelity. AG methods (Li et al., 2021; Zavrtanik et al., 2021; Schlüter et al., 2022; Hu et al., 2024; Gui et al., 2024) only generate anomalies and merge them into the real normal images. NSA (Schlüter et al., 2022) uses Poisson Image Editing (Pérez et al., 2003) to facilitate the fusion of the cropped normal region. However, AG methods require anomaly masks as inputs. If these masks are positioned in an unreasonable manner, the generated images will have low fidelity and consistency. Previous AIG methods (Duan et al., 2023; Zhang et al., 2021a; Niu et al., 2020) are mainly based on GAN. Defect-GAN (Zhang et al., 2021a) cannot generate masks. The masks produced by DFMGAN (Duan et al., 2023) often do not align accurately with anomalies, limiting their utility in training segmentation models. We propose a controllable AIG model based on Stable Diffusion to generate high-fidelity and diverse anomaly images with accurate masks.
Fine-tuning Diffusion Models. Fine-tuning is a potent strategy for enhancing specific capabilities of pre-trained diffusion models (Gal et al., 2022; Zhang et al., 2023; Brooks et al., 2023). Personalized methods (Ruiz et al., 2023; Gal et al., 2022; Chen et al., 2024) utilize a small set of images to fine-tune the diffusion model, thereby generating images of the same object. Several methods for multi-concept image fine-tuning (Kumari et al., 2023; Xiao et al., 2023; Avrahami et al., 2023; Han et al., 2023; Jin et al., 2024) use cross-attention maps to align embeddings with individual concepts in the image. Nevertheless, they do not consider the diversity requirements between different concepts, which is important for industrial anomaly image generation. Thus, we propose a separation and sharing fine-tuning strategy for the different diversity needs of anomalies and products, which independently learns products and anomalies on a shared U-Net.
Mask Prediction with Generation Method. Previous methods on mask prediction for generated images are mainly based on GANs (Zhang et al., 2021b; Li et al., 2022). However, these approaches do not guarantee the generation of accurate masks for exceedingly small datasets. Based on Stable Diffusion (Rombach et al., 2022), some recent methods, i.e., DiffuMask (Wu et al., 2023b), DatasetDM (Wu et al., 2023a) and DatasetDiffusion (Nguyen et al., 2024), produce masks by exploiting the potential of the cross-attention maps. However, due to the low resolution of the cross-attention maps, they are directly interpolated to a higher resolution to match the image size without any auxiliary information, which leads to significant boundary uncertainty. We incorporate the high-resolution features from the VAE decoder as auxiliary information for resolution retrieving, fusing them with the discriminative features of U-Net decoder to generate accurate high-resolution masks.
@@ 3 Method
The training phase of the proposed Separation and Sharing (SeaS) Fine-tuning strategy is shown in Fig. 2. In Sec. 3.1, we introduce the preliminaries of our approach. In Sec. 3.2, we first design an Unbalanced Abnormal Text Prompt, which contains a set of tokens that characterize products and anomalies separately. Subsequently, we propose the Decoupled Anomaly Alignment (DA) loss to bind the anomaly image regions to the anomaly tokens, and leverage Normal-image Alignment (NA) loss to empower the product token to express globally-consistent normal product surface. The two training processes are implemented separately for abnormal and normal images but on a shared U-Net architecture. Then, based on the well-trained U-Net, we design a Refined Mask Prediction branch to generate accurate masks corresponding to the generated anomaly images in Sec. 3.3. Finally, we detail the generation of the abnormal image-mask pairs in Sec. 3.4.
@@ 3.1 Preliminaries
Stable Diffusion. Given an input image $x_{0}$, Stable Diffusion (Rombach et al., 2022) firstly transforms $x_{0}$ into a latent space as $z=\varepsilon(x_{0})$, and then adds a randomly sampled noise $\epsilon\sim N(0,\textbf{I})$ into $z$ as $\hat{z}_{t}=\alpha_{t}z+\beta_{t}\epsilon$, where $t$ is the randomly sampled timestep. Then, the U-Net is employed to predict the noise $\epsilon$. Let $c_{\theta}(\mathcal{P})$ be the CLIP text encoder that maps conditioning text prompt $\mathcal{P}$ into a conditioning vector e. The training loss of Stable Diffusion can be stated as follows:
@@
Cross-Attention Map in U-Net. Aiming to control the generation process, the conditioning mechanism is implemented by calculating cross-attention between the conditioning vector $\textbf{e}\in\mathbb{R}^{Z\times C_{1}}$ and image features $\textbf{v}\in\mathbb{R}^{r\times r\times C_{2}}$ of the U-Net inner layers (Hertz et al., 2022; Chefer et al., 2023; Xie et al., 2023). The cross-attention map $A^{m,l}\in\mathbb{R}^{r\times r\times Z}$ can be calculated as:
@@ 3.2 Separation and Sharing Fine-tuning
Unbalanced Abnormal Text Prompt. Through the experimental observation, we found that the typical text prompt, like a photo of a bottle with defect (Jeong et al., 2023), or damaged bottle (Zhou et al., 2024b), is suboptimal for industrial anomaly generation. The fixed generic semantic words, e.g., damaged, defect, may fail to align with a few training images that contain specific defect types. Therefore, we design the Unbalanced Abnormal (UA) Text Prompt for each anomaly type of each product, i.e.,
@@
where <ob> and <$\texttt{df}_{n}$> ($n\in\{1,2,...,N\}$) are the tokens of the industrial products (short for Normal Token) and the defects (short for Anomaly Token) respectively. As shown in Fig. 3, in SeaS, we separately employ normal images to train the embedding corresponding to <ob>, and abnormal images to train the embeddings corresponding to <$\texttt{df}_{n}$>. Experimental observations indicate that one <ob> is sufficient to express the normal product, while multiple <$\texttt{df}_{n}$> are necessary for controlling the generation of the anomalies. As shown in Fig. 3(a), when we use the UA prompt $\mathcal{P}$ (the dotted green box in (a)), the cross-attention maps in (b) show that different tokens have different responses in the abnormal regions, which indicates that they focus on different attributes of the anomalies, and performing the average operation on the cross-attention maps produces never-seen anomalies. When we use only one <df>, it is difficult to align it to several different anomalies that belong to the same category. Therefore, during inference, if the denoised anomaly feature has a larger distance to <df>, it will be assigned a smaller response by the U-Net, which leads to the “anomaly missing" phenomenon, e.g., the generated images in the case of ($N^{\prime}=1,N=1$). In addition, if we utilize a large number of <$\texttt{df}_{n}$>, we find that each <$\texttt{df}_{n}$> may focus on some local properties of an anomaly, such a case increases the diversity but may reduce the authenticity of the anomalies, as shown in the case $N^{\prime}=1,N=8$. Similarly, if we use multiple learnable <ob>, e.g., $N^{\prime}=4,N=4$, each <ob> pays attention to the local character of the product, which may reduce the authenticity of the product.
@@
where $A_{\text{df}}^{n,l}\in\mathbb{R}^{r\times r\times 1}$ is the cross-attention map corresponding to the $n$-th anomaly token <$\texttt{df}_{n}$>, $N$ is the number of anomaly token in $\mathcal{P}$. $L$ is the total number of U-Net layers used in alignment. $M^{l}$ is the binary mask with $r\times r$ resolution, where the abnormal area is 1 and the background is 0. $A^{l}_{\text{ob}}\in\mathbb{R}^{r\times r\times 1}$ is the cross-attention map corresponding to the normal token <ob>, $\odot$ is the element-wise product. DA loss performs the mandatory decoupling of the anomaly and the product. The first term of DA loss is to align the abnormal area to <$\texttt{df}_{n}$> according to the mask $M^{l}$. The second term of DA loss reduces the response value of $A^{l}_{\text{ob}}$ in the abnormal area, which prevents <ob> from aligning to the abnormal area of $x_{\text{df}}$. Further analysis of how the DA loss ensures the diversity of anomalies is provided in Appendix A.2. Therefore, the total loss for the anomaly image $x_{\text{df}}$ is:
@@
Normal-image Alignment. As we discussed, increasing the number of the normal token <ob> leads to a higher diversity, while may reduce the authenticity of the generated image and destruct global consistency. However, aligning only one <ob> to a few of the training images may suffer from the issue of overfitting. Therefore, we add a Normal-image Alignment (NA) loss to overcome such a dilemma, which is stated as follows,
@@
Instead of aligning the normal region of $x_{\text{df}}$ to <ob>, in calculating the NA loss, we use random noises $\epsilon_{\text{ob}}$ and timesteps $t_{\text{ob}}$ to perform forward diffusion on the normal product images $x_{\text{ob}}$. Then the noisy latent $\hat{z}_{ob}$ and the embedding $\textbf{e}_{\text{ob}}$ corresponding to the normal tokens of $\mathcal{P}$, i.e., “a <ob>", are input into the U-Net in predicting noise, and then calculate the NA loss with $\epsilon_{\text{ob}}$.
Mixed Training. Based on the separated DA loss for abnormal images and NA loss for the normal images, the objective of Separation and Sharing Fine-tuning is expressed as:
@@
In the training process, instead of training a single U-Net model for each anomaly type, we train a unified U-Net model for each product. Specifically, given a product image set, which contains $G$ anomaly categories and some normal images with their corresponding masks. We group all the abnormal images of a product into a unified set $X_{\text{df}}=\{x^{1}_{\text{df}},x^{2}_{\text{df}},..,x^{H}_{\text{df}}\}$. For each anomaly type, we use $\mathcal{P}$ with different sets of anomaly tokens. In addition, we sample a fixed number of normal images to consist of the normal training set $X_{\text{ob}}=\{x^{1}_{\text{ob}},x^{2}_{\text{ob}},..,x^{P}_{\text{ob}}\}$. During each step of our fine-tuning process, we sample same number of images from both $X_{\text{df}}$ and $X_{\text{ob}}$, and mixed them into a batch. We found that such a mixed training strategy not only alleviates the overfitting caused by the limited number of each anomaly type, but also increases the diversity of the anomaly image, while still maintaining reasonable authenticity, as is shown in Fig. 3(c), (i) indicates that the model with mixed training may generate new anomalies, e.g., the anomalies inside the dotted white line. In contrast, the anomalies in (ii) overfit the training images. More ablation studies on the mixed training strategy are shown in Tab. 23 in appendix A.8.

### v2

@@ Abstract
We introduce SeaS, a unified industrial generative model for automatically creating diverse anomalies, authentic normal products, and precise anomaly masks. While extensive research exists, most efforts either focus on specific tasks, i.e., anomalies or normal products only, or require separate models for each anomaly type. Consequently, prior methods either offer limited generative capability or depend on a vast array of anomaly-specific models. We demonstrate that U-Net’s differentiated learning ability captures the distinct visual traits of slightly-varied normal products and diverse anomalies, enabling us to construct a unified model for all tasks. Specifically, we first introduce an Unbalanced Abnormal (UA) Text Prompt, comprising one normal token and multiple anomaly tokens. More importantly, our Decoupled Anomaly Alignment (DA) loss decouples anomaly attributes and binds them to distinct anomaly tokens of UA, enabling SeaS to create unseen anomalies by recombining these attributes. Furthermore, our Normal-image Alignment (NA) loss aligns the normal token to normal patterns, making generated normal products globally consistent and locally varied. Finally, SeaS produces accurate anomaly masks by fusing discriminative U-Net features with high-resolution VAE features. SeaS sets a new benchmark for industrial generation, significantly enhancing downstream applications, with average improvements of $+8.66\%$ pixel-level AP for synthesis-based AD approaches, $+1.10\%$ image-level AP for unsupervised AD methods, and $+12.79\%$ IoU for supervised segmentation models. Code is available at https://github.com/HUST-SLOW/SeaS.
@@ 1 Introduction
In the industrial scenario, generative models are used to synthesise various visual elements, which meet the requirements of different anomaly detection (AD) methods and supervised segmentation models as below.
  • Generating pseudo-anomalies for synthesis-based AD approaches [41, 9].
  • Generating pseudo-normal images for unsupervised AD methods [26, 32, 15].
  • Generating complete anomaly images and corresponding accurate masks for training supervised segmentation models.
@@
Figure 1: (a) AnomalyDiffusion only generates anomalies and edits them onto the input normal images guided by the predefined masks. (b) DFMGAN trains multiple dedicated generators per anomaly type or normal product, and it cannot produce accurate masks. (c) SeaS trains a unified model capable of generating anomaly images and masks for multiple anomaly types, as well as normal images.
These requirements above have been covered by previous algorithms. CutPaste [21] creates anomalies by pasting cropped normal regions onto normal product images (short for normal images). AnomalyDiffusion [17] generates anomalies using diffusion models and edits anomalies onto normal images guided by pre-defined masks (Fig. 1(a)), but it cannot create pseudo-normal images and might suffer from misaligned masks. DFMGAN [11] trains separate models for each normal product and anomaly type (Fig. 1(b)) but cannot produce accurate masks, limiting its effectiveness in training supervised segmentation models. In summary, existing methods either focus only on the generation of normal products or anomalies, or require multiple isolated models to complete all tasks separately. They cannot flexibly use a unified model to tackle them all, i.e., achieving diverse anomalies, authentic normal products, and pixel-accurate masks. In this paper, we study the unified industrial anomaly generation solution, meeting the needs of various downstream tasks.
@@
The novelty of this work stems from a key observation on a single industrial production line: normal products exhibit a globally consistent surface with minor local variations, whereas anomalies exhibit high variability. These characteristics can be effectively captured by U-Net due to its differential learning capability in a diffusion process. Building on this insight, we propose a Separation and Sharing Fine-tuning method (SeaS), using a shared U-Net to model distinct variations. Firstly, to explicitly model the variations of normal products and anomalies, we propose Unbalanced Abnormal (UA) Text Prompt. Its unbalanced design includes one normal token and multiple anomaly tokens, thus decoupling the presentations of the slightly varied normal product surface and diverse anomaly semantics. Secondly, to learn highly-diverse anomalies, we propose a Decoupled Anomaly Alignment (DA) loss to bind the attributes of anomalies to different anomaly tokens of UA. Recombining the decoupled attributes may produce anomalies that have never been seen in the training dataset, therefore increasing the diversity of generated anomalies. Thirdly, for slightly varied normal products, we propose the Normal-image Alignment (NA) loss. It enables the network to learn the key features of the normal product from normal images, so that the normal token of UA expresses the products’ global consistency. The two training processes above are separated but conducted on a shared U-Net. SeaS enables U-Net to simultaneously model the different variations in both normal products and anomalies, representing the discriminative features for mask prediction. However, the low-resolution features of U-Net lead to a coarse mask when predicted directly. Thus, we propose a Refined Mask Prediction (RMP) branch. It combines U-Net features with high-resolution VAE features to generate accurate and crisp masks progressively. The generated anomaly images achieve IS scores by $1.88$ (MVTec AD), $1.27$ (VisA), and $1.95$ (MVTec 3D AD), with IC-LPIPS of $0.34$, $0.26$, and $0.30$. On multiple datasets, SeaS-generated images boost synthesis-based AD approaches by an average $+8.66\%$ pixel-level AP, improve unsupervised AD methods by an average $+1.10\%$ image-level AP, and enhance supervised segmentation models by an average $+12.79\%$ IoU.
@@
  • We propose a unified generative model for industrial visual elements. It achieves diverse anomalies, globally consistent normal products, and pixel-level accurate masks using only one model, which sets a new standard for this field.
  • The newly designed separated and shared fine-tuning models different variations of normal products and anomalies, enabling precise control over their generation, and obtaining discriminative features for mask prediction.
  • SeaS greatly improves the performance of various synthesis-based and unsupervised AD methods, and empowers supervised segmentation models with decent performance.
@@ 2 Related Work
Anomaly Image Generation. Early non-generative methods [10, 21, 41] rely on data augmentation to synthesize pseudo-anomalies, but suffer from low fidelity due to inconsistent anomaly patterns. Some generative methods [34, 17, 13] only generate anomalies and merge them into the real normal images. NSA [34] uses Poisson Image Editing [30] to fuse the cropped normal region. However, these methods cannot create pseudo-normal images and require anomaly masks as inputs, with unreasonable mask positions compromising fidelity and consistency. GAN-based methods [11, 42, 28] generate the entire anomaly images. DFMGAN[11] trains multiple isolated models to generate normal images and anomaly images for each anomaly type, and the produced masks often do not align accurately with anomalies, limiting their utility in training supervised segmentation models. Different from these existing approaches, we propose a unified generative model based on Stable Diffusion to generate diverse anomalies, globally consistent normal products and pixel-level accurate masks.
Fine-tuning Diffusion Models. Fine-tuning is a potent strategy for enhancing specific capabilities of pre-trained diffusion models [12, 43, 6]. Personalized methods [33, 12, 8] utilize a small set of images to fine-tune the diffusion model, thereby generating images of the same object. Several methods for multi-concept image fine-tuning [20, 37, 1, 14, 19] use cross-attention maps to align embeddings with individual concepts in the image. Nevertheless, they do not consider the different variations in different image regions, which is important for industrial anomaly image generation. Thus, we propose a separation and sharing fine-tuning strategy to model the different degrees of variations of anomalies and normal products, which independently learns products and anomalies on a shared U-Net.
Mask Prediction with Generation Method. Previous methods on mask prediction for generated images are mainly based on features in GANs [44, 22]. However, these approaches do not guarantee the generation of accurate masks for exceedingly small datasets. Based on Stable Diffusion [31], some recent methods, i.e., DiffuMask [36], DatasetDM [35] and DatasetDiffusion [27], produce masks by exploiting the potential of the cross-attention maps. However, due to the low resolution of the cross-attention maps, they are directly interpolated to a higher resolution to match the image size without any auxiliary information, which leads to significant boundary uncertainty. We incorporate the high-resolution features from the VAE decoder as auxiliary information for resolution retrieving, fusing them with U-Net decoder features, which are discriminative due to the modelling of the different variations in normal products and anomalies, to generate accurate high-resolution masks.
@@ 3 Method
The training phase of the proposed Separation and Sharing (SeaS) Fine-tuning strategy is shown in Fig. 2. In Sec. 3.1, we introduce the preliminaries of our approach. In Sec. 3.2, we first design an Unbalanced Abnormal Text Prompt, which contains a set of tokens that characterize normal products and anomalies separately. Subsequently, we propose the Decoupled Anomaly Alignment (DA) loss to bind anomaly image regions to anomaly tokens, and leverage Normal-image Alignment (NA) loss to empower normal token to express the globally-consistent normal product surface. The two training processes are implemented separately for abnormal and normal images but on a shared U-Net architecture. Then, based on the well-trained U-Net, we design a Refined Mask Prediction branch to generate accurate masks corresponding to the generated anomaly images in Sec. 3.3. Finally, we detail the generation of abnormal image-mask pairs and normal images in Sec. 3.4.
@@ 3.1 Preliminaries
Stable Diffusion. Given an input image $x_{0}$, Stable Diffusion [31] firstly transforms $x_{0}$ into a latent space as $z=\varepsilon(x_{0})$, and then adds a randomly sampled noise $\epsilon\sim N(0,\textbf{I})$ into $z$ as $\hat{z}_{t}=\alpha_{t}z+\beta_{t}\epsilon$, where $t$ is the randomly sampled timestep. Then, the U-Net is employed to predict the noise $\epsilon$. Let $c_{\theta}(\mathcal{P})$ be the CLIP text encoder that maps conditioning text prompt $\mathcal{P}$ into a conditioning vector e. The training loss of Stable Diffusion can be stated as follows:
@@
Cross-Attention Map in U-Net. Aiming to control the generation process, the conditioning mechanism is implemented by calculating cross-attention between the conditioning vector $\textbf{e}\in\mathbb{R}^{Z\times C_{1}}$ and image features $\textbf{v}\in\mathbb{R}^{r\times r\times C_{2}}$ of the U-Net inner layers [16, 7, 39]. The cross-attention map $A^{m,l}\in\mathbb{R}^{r\times r\times Z}$ can be calculated as:
@@ 3.2 Separation and Sharing Fine-tuning
Unbalanced Abnormal Text Prompt. Through the experimental observation, we found that the typical text prompt, like a photo of a bottle with defect [18], or damaged bottle [46], is suboptimal for industrial anomaly generation. The balanced semantic words for normal products and anomalies may fail to capture their differential variation degrees. Therefore, we design the Unbalanced Abnormal (UA) Text Prompt for each anomaly type of each product, i.e.,
@@
where <ob> and <$\texttt{df}_{n}$> ($n\in\{1,2,...,N\}$) are the tokens of the industrial normal products (short for Normal Token) and the anomalies (short for Anomaly Token) respectively. We use a set of $N$ Anomaly Tokens for each anomaly type, with different sets corresponding to different anomaly types. As shown in Fig. 3, in SeaS, we separately employ normal images to train the embedding corresponding to <ob>, and abnormal images to train the embeddings corresponding to <$\texttt{df}_{n}$>. Experimental observations indicate that one <ob> is sufficient to express normal product, while multiple <$\texttt{df}_{n}$> are necessary for controlling the generation of anomalies. As shown in Fig. 3(a), when we use the UA prompt $\mathcal{P}$ (the dotted green box in (a)), the cross-attention maps in (b) show that different tokens have different responses in the abnormal regions, which indicates that they focus on different attributes of the anomalies, and performing the average operation on the cross-attention maps produces never-seen anomalies. When we use only one <df>, it is difficult to align it to several different anomalies that belong to the same category. Therefore, during inference, if the denoised anomaly feature has a larger distance to <df>, it will be assigned a smaller response by the U-Net, which leads to the “anomaly missing” phenomenon, e.g., the generated images in the case of ($N^{\prime}=1,N=1$). In addition, if we utilize a large number of <$\texttt{df}_{n}$>, we find that each <$\texttt{df}_{n}$> may focus on some local properties of an anomaly, such a case increases the diversity but may reduce the authenticity of the anomalies, as shown in the case $N^{\prime}=1,N=8$. Similarly, if we use multiple learnable <ob>, e.g., $N^{\prime}=4,N=4$, each <ob> pays attention to the local character of the normal product, which may reduce the global consistency of the normal product.
@@
Figure 4: The Refined Mask Prediction (RMP) branch during inference. The Coarse Feature Extraction utilizes features from the up-2 and up-3 layers of the U-Net Decoder to extract coarse features. The cascaded Mask Refinement Module (MRM) further obtains the mask accurately aligned with the anomaly with the assistance of high-resolution features of the VAE Decoder.
@@
where $A_{\text{df}}^{n,l}\in\mathbb{R}^{r\times r\times 1}$ is the cross-attention map corresponding to the $n$-th anomaly token <$\texttt{df}_{n}$>, $N$ is the number of anomaly token in $\mathcal{P}$. $L$ is the total number of U-Net layers used in alignment. $M^{l}$ is the binary mask with $r\times r$ resolution, where the abnormal area is 1 and the background is 0. $A^{l}_{\text{ob}}\in\mathbb{R}^{r\times r\times 1}$ is the cross-attention map corresponding to the normal token <ob>, $\odot$ is the element-wise product. DA loss performs the mandatory decoupling of the anomaly and the normal product. The first term of DA loss is to align the abnormal area to <$\texttt{df}_{n}$> according to the mask $M^{l}$. The second term of DA loss reduces the response value of $A^{l}_{\text{ob}}$ in the abnormal area, which prevents <ob> from aligning to the abnormal area of $x_{\text{df}}$. Further analysis of how the DA loss ensures the diversity of anomalies is provided in Appendix A.2. Therefore, the total loss for the anomaly image $x_{\text{df}}$ is:
@@
Normal-image Alignment. As we discussed, increasing the number of the normal token <ob> leads to a higher diversity, while it may reduce the authenticity of the generated normal image and destruct global consistency. However, aligning only one <ob> to a few of the training images may suffer from the issue of overfitting. Therefore, we add a Normal-image Alignment (NA) loss to overcome such a dilemma, which is stated as follows,
@@
Instead of aligning the normal region of $x_{\text{df}}$ to <ob>, in calculating the NA loss, we use random noises $\epsilon_{\text{ob}}$ and timesteps $t_{\text{ob}}$ to perform forward diffusion on the normal images $x_{\text{ob}}$. Then the noisy latent $\hat{z}_{ob}$ and the embedding $\textbf{e}_{\text{ob}}$ corresponding to the normal tokens of $\mathcal{P}$, i.e., “a <ob>”, are input into the U-Net in predicting noise, and then calculate the NA loss with $\epsilon_{\text{ob}}$.
Mixed Training. Based on the separated DA loss for abnormal images and NA loss for the normal ones, the objective of Separation and Sharing Fine-tuning is formed as:
@@
In the training process, instead of training a single U-Net model for each anomaly type, we train a unified U-Net model for each product. Specifically, given a product image set, which contains $G$ anomaly categories of masked abnormal images and some normal images. We group all the abnormal images of a product into a unified set $X_{\text{df}}=\{x^{1}_{\text{df}},x^{2}_{\text{df}},..,x^{H}_{\text{df}}\}$. For each anomaly type, we use $\mathcal{P}$ with different sets of anomaly tokens. In addition, we sample a fixed number of normal images to consist of the normal training set $X_{\text{ob}}=\{x^{1}_{\text{ob}},x^{2}_{\text{ob}},..,x^{P}_{\text{ob}}\}$. During each step of our fine-tuning process, we sample same number of images from both $X_{\text{df}}$ and $X_{\text{ob}}$, and mixed them into a batch. We found that such a mixed training strategy not only alleviates the overfitting caused by the limited number of each anomaly type, but also increases the diversity of the anomaly image, while still maintaining reasonable authenticity, as shown in Fig. 3(c), (i) indicates that the model with mixed training may generate new anomalies, e.g., the anomalies inside the dotted white line. In contrast, the anomalies in (ii) overfit the training images. More ablation studies on the mixed training strategy are shown in Tab. 12 in appendix A.5.